OpenStudy (anonymous):

find equation of a circle in standard form where 6,-2 and -4,4 are end points of the diameter

OpenStudy (anonymous):

show work please

OpenStudy (anonymous):

Use this equatoin of circle with end of diameter given is (x-x1)(x-x2)+(y-y1)(y-y2)=0

OpenStudy (anonymous):

I dont understand

OpenStudy (precal):

Use graph paper and determine the center of your circle and its radius then sub that information into\[\left( x-h \right)^2+\left( y-k \right)^2=r^2\] (h,k) is the center of the circle and r is the radius